How much 0.113 M NaOHM NaOH solution will completely neutralize 1.25 LL of 0.228 M HCIM HCl solution?


1
Expert's answer
2021-05-11T05:59:11-0400

Solution:

Balanced chemical equation:

NaOH + HCl → NaCl + H2O

According to the equation above: n(NaOH) = n(HCl)

OR: C(HCl) × V(HCl) = C(NaOH) × V(NaOH)

Thus:

Divide to isolate V(NaOH):

V(NaOH) = C(HCl) × V(HCl) / C(NaOH)

Plug in the numbers and solve for V(NaOH):

V(NaOH) = (0.228 M × 1.25 L) / (0.113 M) = 2.52212 L = 2.52 L

V(NaOH) = 2.52 L


Answer: 2.52 L of NaOH solution.